OpenStudy (anonymous):

Solve –x + 6 = x – 8

OpenStudy (anonymous):

We want to get all the x on one side and all the numbers on the other side. What do you think we should do first.

jhonyy9 (jhonyy9):

add to both sides x and after this add to both sides 8 hence will get -x+x+6+8=x+x-8+8 14=2x x=14/3 =7 x=7